OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. @iGreen.

OpenStudy (anonymous):

Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (anonymous):

We're gonna write it like: \(f(x) = p(1 + r)^x\) So figure out what you want the population to start at and what the rate of growth is.

OpenStudy (anonymous):

for creating the exponential fraction, would i use f(x)=P(1+r)^x or f(x)=a(b)^x

OpenStudy (anonymous):

ssorry internet connection went out

OpenStudy (anonymous):

do i want to chose what i want the population to start at?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Yes, and the rate

OpenStudy (anonymous):

so is this fine principal = 123 rate = 14% ??

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Yes..so we have: \(f(x) = 123(1.14)^x\)

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

now it says A local snail population grows according to the function g(x) = 200(1.03)2^x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. how do i do the first one? it never showed me how in the lesson or previous lessons

OpenStudy (anonymous):

would the key features be the principal and the rate of growth too?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@sleepyjess